OpenStudy (anonymous):

Determine the interval of convergence for the power series: summation n=0 to infinity (x-1)^n/ln n. Please help as I don't even know where to start!

OpenStudy (anonymous):

ratio test for this one

OpenStudy (anonymous):

it is pretty clearly 1 from your eyeballs, but we can take all the steps if you like

OpenStudy (anonymous):

btw i hope the sum starts at \(n=1\) not \(n=0\) since \(\ln(0)\) is undefined

OpenStudy (anonymous):

you start by using the ratio test \[\vert\frac{a_{n+1}}{a_n}\vert\]

OpenStudy (anonymous):

in this case it is \[\\\frac{(x-1)^{n+1}}{\ln(n+1)}\times \frac{\ln(n)}{(x-1)^n}\] in absolute value took me a minute to type it with me so far?

OpenStudy (anonymous):

Yes. I started by using the Ratio Test and ended up with the absolute value of x-1

OpenStudy (anonymous):

it does start at n=0 by the way

OpenStudy (anonymous):

yes, that is right as \[\lim_{n\to \infty}\frac{\ln(n+1)}{\ln(n)}=1\]you get \[|x-1|\] as your limit this will converge if the limit is less than \(1\) so the radius of convergence is 1, and the interval of convergence is what you get when you solve \[|x-1|<1\] for \(x\)

OpenStudy (anonymous):

if it starts at \(n=0\) then either it is a trick question, or more likely just a typo

OpenStudy (anonymous):

Knowing my professor, it's a typo :)

OpenStudy (anonymous):

lol clear how to solve \[|x-1|<1\] right? you do it in your head

OpenStudy (anonymous):

not done yet however you have to check the endpoints of the interval

OpenStudy (anonymous):

I don't understand how to go further. I feel like the interval will be -1<x<1 but I don't understand how to go about finding and checking the endpoints.

OpenStudy (anonymous):

hold the phone

OpenStudy (anonymous):

the inequality you get is \[|x-1|<1\] solve like in elementary algebra \[|x-1|<1\\ -1<x-1<1\\ 0<x<2\]

OpenStudy (anonymous):

Lord I haven't had algebra since high school, which was 15 years ago! So this doesn't even look familiar to me

OpenStudy (anonymous):

hey @satellite73

OpenStudy (anonymous):

better gut used to it i guess \(|x-1|<1\) says the distance between \(x\) and \(1\) is less than \(1\) that is why i thought you can do it in your head 

OpenStudy (anonymous):

its demi remember you helped me with many questions thank you once again

OpenStudy (anonymous):

oh thats ok. not right now but i will when i do

OpenStudy (anonymous):

so we have now only to check the endpoints of the interval i.e. replace \(x\) by \(0\) and then replace \(x\) by \(2\) and see what you get for the sum

OpenStudy (anonymous):

with me so far?

OpenStudy (anonymous):

if you put \(x=0\) you get \(0-1=-1\) so the sum becomes \[\sum\frac{(-1)^n}{\ln(n)}\]

OpenStudy (anonymous):

replace it in the inequality x-1 < 1?

OpenStudy (anonymous):

lets back up a second to make sure it is clear how to solve these inequalities

OpenStudy (anonymous):

If I seem a little slow on the uptake it's because I have a professor that skips from step A to step D without showing how he got there

OpenStudy (anonymous):

thanks for dumbing it down for me

OpenStudy (anonymous):

\[|x-a|<p\] says the distance between \(x\) and \(a\) is less than \(p\) you solve it by starting with \[-p<x-a<p\] then add \(a\) to both sides and get \[-p+a<x<p+a\]

OpenStudy (anonymous):

for example if you have to solve \[|x-2|<3\]start with \[-3<x-2<3\]add \(2\) and get \[-1<x<5\]

OpenStudy (anonymous):

oh. i see now

OpenStudy (anonymous):

if this does not ring a bell, and i am guessing it doesn't, you should review it because you will be solving lots of them

OpenStudy (anonymous):

ok so now we got to the point where we know the limit is \(|x-1|\) and we have to make sure that is is less than 1 so we solve \[|x-1|<1\] and arrive at \(0<x<2\)

OpenStudy (anonymous):

well my final exam is next week and there will be one question about intervals of convergence so i'm hoping i can just get throught it.....know what i mean?

OpenStudy (anonymous):

ok. with you so far

OpenStudy (anonymous):

yeah, but that part is not really that hard in defence of your typo making prof, he or she probably just thought that part was obvious now we check the endpoints of the interval replace \(x\) by \(0\)

OpenStudy (anonymous):

when you replace \(x\) by \(0\) you get \[\sum\frac{(-1)^n}{\ln(n)}\] is that part clear?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it clear that because the terms alternate in sign, and the terms go to zero, that this sum converges? that is all you have to check if it alternate do the terms go to zero or not

OpenStudy (anonymous):

when you say alternate in sign do you mean that it's (-1)^n when you plug in 0 and (1)^n when you plug in 2?

OpenStudy (anonymous):

yes, that is what i mean \((-1)^n\) is either \(-1\) or \(1\) depending on whether \(n\) is odd or even

OpenStudy (anonymous):

ooh no no that is not what i mean let's back up again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\sum\frac{(-1)^n}{\ln(n)}\]is what you get when you replace \(x\) by \(0\)

OpenStudy (anonymous):

THAT series alternates in sign because of the \((-1)^n\) part i.e. \((-1)^n\) alternates in sign it is either \(-1\) or it is \(1\) depending, as i said, if \(n\) is odd or if \(n\) is even

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now if you replace \(x\) by \(2\) you get what you said \[\sum\frac{1}{\ln(n)}\]

OpenStudy (anonymous):

noting of course that \((1)^n=1\) no matter what \(n\) is

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now the question is does \[\sum\frac{1}{\ln(n)}\]converge what do you think?

OpenStudy (anonymous):

Honestly, I have no idea. I really haven't understood anything we've been doing in class the past month. And convergence def confuses me.

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @satellite73 yes, that is right as \[\lim_{n\to \infty}\frac{\ln(n+1)}{\ln(n)}=1\]you get \[|x-1|\] as your limit \(\color{blue}{\text{End of Quote}}\) Shouldn't the limit be 0?

OpenStudy (anonymous):

then you have some catching up to do

OpenStudy (anonymous):

@SithsAndGiggles isn't the limit 1?

OpenStudy (anonymous):

Oh my mistake, I read that as \(\ln(n+1)-\ln n\) for some reason... Carry on!

OpenStudy (anonymous):

in any case the log grows very very very slowly not fast enough for the sum to converge

OpenStudy (anonymous):

@SithsAndGiggles same thing, same limit right?

OpenStudy (anonymous):

Well, \(\ln(n+1)-\ln n=\ln\dfrac{n+1}{n}\), which tends to 0, but \(\dfrac{\ln(n+1)}{\ln n}\) tends to 1.

OpenStudy (anonymous):

yes, of course

OpenStudy (anonymous):

so no, not the same thing, not the same limit

OpenStudy (anonymous):

so they both diverge?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the first one converges because it alternates and the terms go to zero the second one is divergent because it does not alternate, and the terms do not go to zero fast enough

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so finally we see that the "interval of convergence" is \[[0,2)\] that took a while, hope something became clear

OpenStudy (anonymous):

I know. I'm sorry! I really appreciate the time you spent explaining it to me.

OpenStudy (anonymous):

you were a tremendous help

OpenStudy (anonymous):

you might want to look and see what kind of series converge, and which series diverge because even if there is only one "interval of convergence" question on the test there might be a question with just series for example \[\sum\frac{n}{n^2+2}\] do you think that converges?

OpenStudy (anonymous):

oh no don't be sorry, if i hadn't wanted to help i wouldn't have yw

OpenStudy (anonymous):

maybe it converges based on the comparison test???

OpenStudy (anonymous):

no it diverges here is the easy reason the numerator is a polynomial of degree 1 the denominator is a polynomial of degree 2 since the difference in degrees is 1 it does NOT converge the degree of the denominator has to be larger than the degree of the numerator by MORE than 1

OpenStudy (anonymous):

oh yes. i remember that now

OpenStudy (anonymous):

you could compare it to the well known divergent series \[\sum\frac{1}{n}\] if you wanted an actual test

OpenStudy (anonymous):

if he or she is tricky, maybe a problem like \[\sum\frac{1}{n\ln(n)}\]where the denominator is not a polynomial

OpenStudy (anonymous):

thanks for the advise. will def look over examples

OpenStudy (anonymous):

yw btw integral test shows the previous one diverges as well

OpenStudy (anonymous):

Ok. Awesome help!